OpenStudy (anonymous):

f(x) = 5x + 3 g(x) = 2x^2-2x + 2 Find (fg(x) )

OpenStudy (anonymous):

so @jim_thompson5910 for this one which one should i start out with?

jimthompson5910 (jim_thompson5910):

there's no circle I'm assuming, so it looks like you're multiplying these functions

jimthompson5910 (jim_thompson5910):

so (f * g)(x) = f(x) * g(x) (f * g)(x) = (5x+3) * (2x^2 - 2x + 2)

jimthompson5910 (jim_thompson5910):

expand and simplify

OpenStudy (anonymous):

uh one sec

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

10x^3-4x^2+4x+6

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

yay ty!!!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

yw